Q.1 A producer faces a fixed cost of Rs. 50 and a variable cost


Rs. 5 per unit of output when he produces less than 200 units of


output.


• Assuming that total cost function is linear, determine the


equation of the total cost of function.


• What is the break-even level of output if the price of output is


Rs. 10 per unit. Also draw the relevant graphs.


• Determine the producer’s net revenue if output is 12 units and


the price of output is Rs. 5, Rs. 10 and Rs. 15 respectively.

1
Expert's answer
2021-12-23T12:03:27-0500

TC=FC+VC

FC=Rs50+5Q

"BEP=FC\/Contribution Margin"

=50/(10-5)

50/5

=10





TR=P.Q

i) TR=Rs5*12= Rs 60

ii) TR=Rs10*12=Rs120

iii) TR=Rs15*12=Rs180